# citations.bib

### v1

@@ Abstract
Detecting vulnerabilities within compiled binaries is challenging due to lost high-level code structures and other factors such as architectural dependencies, compilers, and optimization options. To address these obstacles, this research explores vulnerability detection by using natural language processing (NLP) embedding techniques with word2vec, BERT, and RoBERTa to learn semantics from intermediate representation (LLVM) code. Long short-term memory (LSTM) neural networks were trained on embeddings from encoders created using approximately 118k LLVM functions from the Juliet dataset. This study is pioneering in its comparison of word2vec models with multiple bidirectional transformer (BERT, RoBERTa) embeddings built using LLVM code to train neural networks to detect vulnerabilities in compiled binaries. word2vec Continuous Bag of Words (CBOW) models achieved 92.3% validation accuracy in detecting vulnerabilities, outperforming word2vec Skip-Gram, BERT, and RoBERTa. This suggests that complex contextual NLP embeddings may not provide advantages over simpler word2vec models for this task when a limited number (e.g. 118K) of data samples are used to train the bidirectional transformer-based models. The comparative results provide novel insights into selecting optimal embeddings for learning compiler-independent semantic code representations to advance machine learning detection of vulnerabilities in compiled binaries.
@@ I Introduction
Microsoft is one of the dominant players within the server and desktop operating system market; therefore, vulnerabilities in its software can negatively impact millions of systems around the world. Some of the most prolific worms in the world have exploited vulnerabilities in software developed by Microsoft; these worms include Code Red [RN72], which exploited a vulnerability in Microsoft IIS, Sasser [RN73], which exploited the Windows LSASS process, and Conficker [RN74] and WannaCry [RN75] which exploited SMB vulnerabilities in Windows systems. The financial impact of recovering from these worms is estimated to be billions of dollars (USD) [RN76]. Microsoft’s closed-source nature  [RN77] prevents organizations from reviewing code written in a high-level programming language for vulnerabilities before deploying it within their environments. Microsoft is a single example of an organization that produces closed-source products, but is certainly not the only company. Most commercial organizations protect their intellectual property by not releasing their source code to the public. Therefore, organizations that use these commercial products rely entirely on the organizations that create the software to follow secure coding practices. However, they cannot independently verify the security of the code.
Locating vulnerabilities within compiled binaries can be difficult, because significant information is lost when source code is compiled on a target platform. The names of variables, programming structures, and code comments are reduced to low-level machine code. Furthermore, machine code is significantly different for each platform on which it is compiled (e.g., x86/x64, MIPS, SPARC, ARM, etc.). To solve these issues, many researchers have turned to machine learning in the hope that it may provide a solution for locating vulnerabilities in compiled binaries.
@@
Although researchers have explored the use of specific natural language processing (NLP) encoding techniques in the context of decompiled code, research has not yet focused on comparative analysis using the most popular NLP encoding models to identify which embedding techniques perform best in the task of identifying vulnerabilities in compiled binaries. Very few studies have used bidirectional transformer-based models to acquire semantics within decompiled code for this task. RoBERTa was used by [RN67] to learn semantics within the x86 assembly language, and [gallagherllvm] used this model to learn the semantics of the LLVM code. Additionally, [koo2021semantic] uses BERT for code similarity detection, and [han2022binary] trained a BERT model using pcode and used the embeddings to differentiate between vulnerable and non-vulnerable code. Besides [gallagherllvm] and [han2022binary], little to no research has explored the use of unidirectional encoders to identify vulnerabilities in compiled binaries that have been lifted to languages such as LLVM. Furthermore, according to our research, there have been no comparative studies on the performance of neural networks trained using embeddings from common NLP models (word2vec, BERT, and RoBERTa) to differentiate between vulnerable and non-vulnerable LLVM code. This paper explores the use of multiple bidirectional encoding models (BERT & RoBERTa) to evaluate their effectiveness in generating embeddings that can be used to accurately detect vulnerabilities in compiled binaries. The models built using these bidirectional embedding models are compared to those generated using word2vec: CBOW and Skip-Gram embeddings.
This paper can be viewed as an extension of previous work of [RN9] and [schaad2022deeplearningbased]. In these studies, the researchers compiled code samples from the SARD dataset, lifted them to LLVM using the RetDec tool, and used the lifted code to build a word2vec model. Once this model was built, the embeddings from the word2vec model were provided to neural networks to train them to differentiate between vulnerable and non-vulnerable code. However, the research articulated in this paper acknowledges the limitations of word2vec’s ability to create context-aware embeddings and explores the alternative use of bidirectional transformer-based models.
@@ II Design and Implementation
The current research uses machine learning techniques to identify vulnerabilities in compiled binaries by first using NLP bidirectional encoders to learn code-level semantics in lifted LLVM code. The embeddings generated from these encoders were then used to train LSTM neural networks, and the performance of these neural networks was compared with neural networks trained using simpler word2vec embedding models. An overview of the steps involved in this process is as follows.
@@
    All code samples within the NIST SARD [RN11] Juliet dataset were compiled into object files, and the object files were lifted to LLVM. Although the Juliet dataset contains several classes of vulnerabilities and a mix of code samples containing vulnerable and non-vulnerable functions, the primary purpose of the dataset was to learn the semantics between tokenized LLVM instructions. Therefore, tracking vulnerability classes and vulnerable/non-vulnerable status was not required. This process is located in II-A.
@@ II-A Embedding Dataset Creation
The dataset chosen for this study is the C/C++ code samples from the Juliet 1.3 dataset [RN89] within the larger SARD dataset [RN11]. The rationale for selecting this dataset lies in its use in several other research projects [RN50, RN43, RN67, RN56, RN47, RN46, RN55, schaad2022deeplearningbased]. Choosing a dataset used within multiple research projects provides a basis for a meaningful comparison with the previous corpus of knowledge, as well as for easier reproducibility of results.
@@
The Juliet data set, which consists of 101,231 C/C++ files, was compiled to separate object files on a Kali 2023.3 and Windows 10 system. Once each source code sample was compiled into an object file, the RetDec tool was used to convert the object files into an LLVM file. Many articles analyzed during the literature review described research on building models to detect vulnerabilities within x86/x64 assembly language instructions. However, this approach limits the applicability of the results to that instruction set, and the same process cannot be applied to additional architectures. Thus, the current research project used RetDec [RN22] to lift binaries from x86/x64 to an intermediary language called LLVM. The compiler-agnostic [RN18] nature of LLVM allows for future expansion of datasets for samples compiled using different architectures. Additionally, the use of LLVM should help reduce some differences between compiler-specific optimizations [RN18]. Theoretically, a lifted binary for x86/x64 should be identical to a lifted binary compiled for a different architecture (e.g., ARM, SPARK, etc.).
@@
TABLE I: Embedding Training Dataset
Operating System Object Files LLVM Files
Extracted
Functions
Windows 10 96,061 94,225
Kali Linux
2023.3
65,972 65,972
Total
(Pre Duplicate
Removal)
162,033 160,197 2,479,856
Total
(Post Duplicate
Removal)
117,742
@@ II-B Neural Network Dataset Creation
Creating the second dataset involved isolating the Stack-Based Buffer Overflow (CWE-121) LLVM samples from the dataset used to train the embedding models. The breakdown of the number of functions within this dataset is in Table  II.
TABLE II: LSTM Training Dataset
Operating System Object Files LLVM Files
Extracted
Functions
Windows 10 7,871 7,009
Kali Linux
2023.3
4,949 4,949
Total
(Pre Duplicate
Removal)
12,820 11,958
31,301
-17,501 (Clean)
-13,800 (Vuln)
Total
(Post Duplicate
Removal)
6,888
-4,319 (Clean)
-2,569 (Vuln)
Total
(Removal of
Functions More
than 2048 Tokens)
6,773
-4,227 (Clean)
-2,546 (Vuln)
@@ II-C LLVM Preprocessing
Once the LLVM datasets were generated, a sampling of the LLVM files was manually analyzed to determine the best processing approach. The primary goal of the embedding algorithms was to learn the semantics of the LLVM code. Thus, preprocessing was performed to reduce the number of tokens to focus more on execution order and less on function-specific code offsets. A Jupyter Notebook was created that performs the actions listed in Table  III.
TABLE III: Preprocessing Steps
@@
labels, and stack
pointers were renamed
@@
Variables surrounded
by quotes were
converted to unique
tokens used to track
order of appearance
@“misc_var” “stringvar_x”
Numeric constants
greater than nine
were given a unique
identifier based on
order of appearance
12345
-12345
instconstpos_x
instconstneg_x
@@
opencurlybrackets
dotpuncmark
commapuncmark
Note: x is a number used to track the order of appearance in the function.
@@
The first of these techniques, CBOW, generates a vector that uniquely represents a particular word using the words before and after. It is especially good at identifying similarities between words because similar words are often found within the same word patterns. For example, in the sentence “one of my favorite foods to eat is ice cream.”, the word ice could easily be replaced with whipped or even sour. Thus, these words would have vectors that highlight that they are similar. However, the word road would have a vector that is quite different than ice, whipped, or sour because the words before and after the word are generally quite different from those for the other words. The CBOW technique predicts a missing word using the words surrounding the target word. The second technique used by word2vec, Skip-Gram, attempts to predict surrounding words using a target word, unlike CBOW, which predicts a target word based on the surrounding words.
The word2vec: CBOW and word2vec: Skip-Gram models were created using the gensim Python library. Each preprocessed LLVM function was tokenized using the word tokenizer function of the nltk library. The model was built using all the 117,742 samples from the embedding model dataset. There were 5,224 unique tokens used to train these models, but the vector size of the output embedding was set to one hundred. The window size was set to five and the minimum count required for each token was set to one.
@@ II-E BERT and RoBERTa Model Creation
The code for the BERT model was released in 2019 [devlin2019bert] and stands for Bidirectional Encoder Representations from Transformers. BERT is a bidirectional encoder that learns token embeddings by masking out tokens within the input data and attempting to predict the missing token. Whereas some of the current models at the time were unidirectional, BERT proposed a bidirectional model that scored substantially higher on various NLP tasks than its predecessors. RoBERTa (Robustly Optimized BERT pre-training Approach) was also released in 2019 [liu2019roberta]. As its name suggests, RoBERTa is similar to BERT, with some optimizations used to increase its performance.
The custom BERT model used in the current research for learning semantics between LLVM instructions was created using the HuggingFace BertForMaskedLM library, and the BertWordPieceTokenizer was used to convert LLVM functions to tokens. Similarly, the RoBERTa model was generated using the RobertaForMaskedLM library and the ByteLevelBPETokenizer library. In both cases, 12 hidden layers with 100 nodes were used, each with an output layer of 100. The maximum vocabulary length was 2048, and batch sizes were limited to one due to GPU memory limitations. Training the BERT and RoBERTa models requires a training and validation dataset. Thus, the 117,742 samples from the embedding model dataset were split into two datasets; the training dataset comprised 90% of the overall dataset (i.e., 105,967) and the validation dataset was 10% of the dataset (i.e., 11,775).
For both the BERT and RoBERTa embedding models, the loss of training and validation steps was shown every 1,000 steps. The best loss in the validation set for the BERT model was achieved after step 42,000, and the best loss for the RoBERTa model was achieved at step 41,000 as shown in Tables  IV and  V. These were the embedding models used to create the embeddings provided to the LSTM models.
TABLE IV: BERT Five Best Loss Scores by Validation Step
Step Training Set Validation Step
42,000 2.698300 2.639552
41,000 2.694500 2.639749
40,000 2.689700 2.642417
38,000 2.701500 2.646453
39,000 2.690200 2.647793
TABLE V: RoBERTa Five Best Loss Scores by Validation Step
Step Training Set Validation Step
41,000 2.733500 2.701580
42,000 2.741900 2.703620
39,000 2.736500 2.708259
40,000 2.737100 2.709357
37,000 2.763900 2.712307
@@ II-F LSTM Neural Network Configuration
For the current research project, the LSTM neural networks comprised two hidden layers with 128 neurons each and 50 epochs to train each network. The Leaky Rectified Linear Unit (LeakyReLU) activation function was chosen because recurrent neural networks are prone to the problem of vanishing or exploding gradients [RN25, p. 335], and the LeakyReLU activation function helps avoid this problem. Each layer of the LSTM model was created using the TensorFlow and Keras libraries. A dropout rate of 20% was used between the LSTM layers and a single dense layer of a single neuron was used in the output layer. Furthermore, the layers of each embedding model were frozen so they could not be updated as the LSTM was being trained to differentiate between vulnerable and non-vulnerable code. Five LSTMs were generated for each embedding model. The first LSTM used the Adam optimizer with an initial learning rate of 0.0001. The second LSTM used the Stochastic Gradient Descent (SGD) optimizer with an initial learning rate and momentum of 0.01. Finally, the remaining three LSTM neural networks used the SGD optimizer with an initial learning rate of 0.0001 and momentums of 0.01, 0.001, and 0.0001. The dataset was randomized and divided into two subsets. The first subset was used to train each neural network and consisted of 80% (i.e., 5,418 samples) of the overall dataset; the second subset was used for model validation and consisted of 20% (i.e., 1,355) of the total dataset. The same training and validation subsets were used for each neural network.
III Results
Accuracy, precision, recall, and F1 score were evaluated for each LSTM neural network created from the embeddings generated from each embedding model.
III-A word2vec: CBOW
The metrics in Table  VI were recorded when each model achieved its highest accuracy. Although the highest accuracy, 92.3%, was obtained when the Adam optimizer was used with a learning rate of 0.0001, the model built using the SGD optimizer model with a learning rate and momentum of 0.01 was a close second with an accuracy of 91.7%. LSTMs created using the SGD optimizer with learning rates of 0.0001 had gradients that tended to stall in the middle epochs but started to spike in accuracy toward the latter epochs. However, in these final epochs, the model’s accuracy improved rapidly until the final epoch. Thus, it is likely that more epochs would have resulted in greater accuracy.
TABLE VI: word2vec: CBOW Metrics
Hyper-
parameters
Epoch Loss Accuracy Precision Recall
F1-
Score
Adam
-LR: 0.0001
50 0.2216 92.3% 87.9% 91.9% 89.9%
SGD
-LR: 0.01
-Mom: 0.01
49 0.2129 91.7% 88.2% 89.8% 89.0%
SGD
-LR: 0.0001
-Mom: 0.0001
50 0.4490 77.1% 66.5% 78.1% 71.9%
SGD
-LR: 0.0001
-Mom: 0.001
50 0.5164 72.2% 65.5% 54.5% 59.5%
SGD
-LR: 0.0001
-Mom: 0.01
50 0.5289 71.0% 65.1% 47.8% 55.2%
III-B word2vec: Skip-Gram
@@ III Results
The metrics in Table  VII were calculated when the accuracy of the validation set was the highest for each model. The results for the SGD optimizer with a learning rate of 0.0001 and a momentum of 0.01, 0.001, and 0.0001 have been excluded because the model struggled to achieve accuracy gains beyond the base accuracy of placing all samples in a single category. The highest performing model built using word2vec Skip-Gram embeddings attained an accuracy of 90.6%.
TABLE VII: word2vec: Skip-Gram Metrics
Hyper-
parameters
@@
Epoch Loss Accuracy Precision Recall
F1-
Score
Adam
-LR: 0.0001
40 0.2201 90.6% 87.4% 87.6% 87.5%
SGD
-LR: 0.01
-Mom: 0.01
30 0.3143 83.8% 74.9% 85.6% 79.9%
III-C BERT
The LSTM model that performed best with BERT embeddings was built using the Adam optimizer with a learning rate of 0.0001. Table  VIII highlights that the LSTM models trained using BERT embeddings hovered around an accuracy of 81%, with the best-performing model achieving an accuracy of 84.1%.
TABLE VIII: BERT Metrics
Hyper-
parameters
@@ III-C BERT
Epoch Loss Accuracy Precision Recall
F1-
Score
Adam
-LR: 0.0001
35 0.2985 84.1% 80.2% 76.4% 78.2%
SGD
-LR: 0.0001
-Mom: 0.001
42 0.3828 83.0% 74.3% 83.7% 78.7%
SGD
-LR: 0.01
-Mom: 0.01
16 0.3713 81.8% 70.0% 89.8% 78.7%
SGD
-LR: 0.0001
-Mom: 0.01
37 0.4003 81.3% 72.4% 81.1% 76.5%
SGD
-LR: 0.0001
-Mom: 0.0001
48 0.3975 81.0% 72.6% 78.9% 75.7%
III-D RoBERTa

### v2

@@ Abstract
Detecting vulnerabilities within compiled binaries is challenging due to lost high-level code structures and other factors such as architectural dependencies, compilers, and optimization options. To address these obstacles, this research explores vulnerability detection using natural language processing (NLP) embedding techniques with word2vec, BERT, and RoBERTa to learn semantics from intermediate representation (LLVM IR) code. Long short-term memory (LSTM) neural networks were trained on embeddings from encoders created using approximately 48k LLVM functions from the Juliet dataset. This study is pioneering in its comparison of word2vec models with multiple bidirectional transformer (BERT, RoBERTa) embeddings built using LLVM code to train neural networks to detect vulnerabilities in compiled binaries. word2vec Skip-Gram models achieved 92% validation accuracy in detecting vulnerabilities, outperforming word2vec Continuous Bag of Words (CBOW), BERT, and RoBERTa. This suggests that complex contextual embeddings may not provide advantages over simpler word2vec models for this task when a limited number (e.g. 48K) of data samples are used to train the bidirectional transformer-based models. The comparative results provide novel insights into selecting optimal embeddings for learning compiler-independent semantic code representations to advance machine learning detection of vulnerabilities in compiled binaries.
@@ I Introduction
Microsoft is one of the dominant players within the server and desktop operating system market; therefore, vulnerabilities in its software can negatively impact millions of systems around the world. Some of the most prolific worms in the world have exploited vulnerabilities in software developed by Microsoft; these worms include Code Red [RN72], which exploited a vulnerability in Microsoft IIS, Sasser [RN73], which exploited the Windows LSASS process, and Conficker [RN74] and WannaCry [RN75] which exploited SMB vulnerabilities in Windows systems. The financial impact of recovering from these worms is estimated to be billions of dollars (USD) [RN76]. Microsoft’s closed-source nature  [RN77] prevents organizations from reviewing code written in a high-level programming language for vulnerabilities before deploying it within their environments. Microsoft is a single example of an organization that produces closed-source products but is certainly not the only company. Most commercial organizations protect their intellectual property by not releasing their source code to the public. Therefore, organizations that use these commercial products rely entirely on the organizations that create the software to follow secure coding practices. However, they cannot independently verify the security of the code.
Locating vulnerabilities within compiled binaries can be difficult because significant information is lost when source code is compiled on a target platform. The names of variables, programming structures, and code comments are reduced to low-level machine code. Furthermore, machine code is significantly different for each platform on which it is compiled (e.g., x86/x64, MIPS, SPARC, ARM, etc.). To solve these issues, many researchers have turned to machine learning in the hope that it may provide a solution for locating vulnerabilities in compiled binaries.
@@
Although researchers have explored the use of specific NLP encoding techniques in the context of decompiled code, research has not yet focused on comparative analysis using the most popular NLP encoding models to identify which embedding techniques perform best in the task of identifying vulnerabilities in compiled binaries. Very few studies have used bidirectional transformer-based models to acquire semantics within decompiled code for this task. RoBERTa was used by [RN67] to learn semantics within the x86 assembly language, and [gallagherllvm] used this model to learn the semantics of the LLVM code. Additionally, [koo2021semantic] uses BERT for code similarity detection, and [han2022binary] trained a BERT model using pcode and used the embeddings to differentiate between vulnerable and non-vulnerable code. Besides [gallagherllvm] and [han2022binary], little to no research has explored the use of unidirectional encoders to identify vulnerabilities in compiled binaries that have been lifted to languages such as LLVM. Furthermore, according to our research, there have been no comparative studies on the performance of neural networks trained using embeddings from common NLP models (word2vec, BERT, and RoBERTa) to differentiate between vulnerable and non-vulnerable LLVM code. This paper explores the use of multiple bidirectional encoding models (BERT & RoBERTa) to evaluate their effectiveness in generating embeddings that can be used to accurately detect vulnerabilities in compiled binaries. The models built using these bidirectional embedding models are compared to those generated using word2vec: CBOW and Skip-Gram embeddings.
This paper can be viewed as an extension of previous work of [RN9] and [schaad2022deeplearningbased]. In these studies, the researchers compiled code samples from the SARD dataset, lifted them to LLVM using the RetDec [RN22] tool, and used the lifted code to build a word2vec model. Once this model was built, the embeddings from the word2vec model were provided to neural networks to train them to differentiate between vulnerable and non-vulnerable code. However, the research articulated in this paper acknowledges the limitations of word2vec’s ability to create context-aware embeddings and explores the alternative use of bidirectional transformer-based models.
@@ II Design and Implementation
The current research uses machine learning techniques to identify vulnerabilities in compiled binaries by first using NLP bidirectional encoders to learn code-level semantics in lifted LLVM code. The embeddings generated from these encoders were then used to train LSTM neural networks, and the performance of these neural networks was compared with neural networks trained using simpler word2vec embedding models. An overview of the steps involved in this process is as follows. Fig. 1 is a visual representation of steps one through four, and 2 depicts steps five and six of this process.
@@
    All code samples, with the exception of samples containing CWE-121, in the NIST SARD Juliet [RN89] dataset were compiled into object files, and the object files were lifted to LLVM. Although the Juliet dataset contains several classes of vulnerabilities and a mix of code samples containing vulnerable and non-vulnerable functions, the primary purpose of the dataset was to learn the semantics between tokenized LLVM instructions. Therefore, tracking vulnerability classes and vulnerable/non-vulnerable status was not required. This process is located in II-A.
@@
Figure 1: Building Embedding Models Using Lifted Code
@@ II-A Embedding Dataset Creation
The dataset chosen for this study is the C/C++ code samples from the Juliet 1.3 dataset within the larger SARD dataset [RN11]. The rationale for selecting this dataset lies in its use in several other research projects [RN50, RN43, RN67, RN56, RN47, RN46, RN55, schaad2022deeplearningbased]. Choosing a dataset used within multiple research projects provides a basis for a meaningful comparison with the previous corpus of knowledge, as well as for easier reproducibility of results.
@@
The Juliet data set, excluding samples containing CWE-121, was compiled to separate object files on a Kali 2023.3 system [^1] The samples were compiled into object files on a Windows system using both Clang and CL. However, RetDec removed calls to the C Standard Library when lifting the object files to LLVM. As a result, these nonrepresentative samples were excluded from the datasets used to train LLMs and LSTM models.. Once each source code sample was compiled into an object file, the RetDec tool converted the object files into an LLVM file. Many articles analyzed during the literature review described research on building models to detect vulnerabilities within x86/x64 assembly language instructions. However, this approach limits the applicability of the results to that instruction set, and the same process cannot be applied to additional architectures. Thus, the current research project used RetDec to lift binaries from x86/x64 to an intermediary language called LLVM. The compiler-agnostic [RN18] nature of LLVM allows for future expansion of datasets for samples compiled using different architectures. Additionally, using LLVM should help reduce some differences between compiler-specific optimizations [RN18]. Theoretically, a lifted binary for x86/x64 should be identical to a lifted binary compiled for a different architecture (e.g., ARM, SPARK, etc.).
@@ II-B Neural Network Dataset Creation
Creating the second dataset involved compiling Juliet samples containing stack-based buffer overflows (CWE-121) into object files and using RetDec to lift these samples to LLVM. Once lifted to LLVM, the clean and vulnerable functions were extracted from the LLVM files. The breakdown of the number of functions within this dataset is in Table  I.
TABLE I: Embedding and LSTM Training Datasets
Training (size)
Dataset Purpose LLM LSTM
Object Files 61,023 4,949
LLVM Files 61,023 4,949
Extracted Functions:
12,069   Clean: 7,011 Vuln: 5,058
3,901   Clean: 2,452 Vuln: 1,449
3,802   Clean: 2,386 Vuln: 1,416
@@ II-C LLVM Preprocessing
Once the LLVM datasets were generated, a sampling of the LLVM files was manually analyzed to determine the best processing approach. The primary goal of the embedding algorithms was to learn the semantics of the LLVM code. Thus, preprocessing was performed to reduce the number of tokens to focus more on execution order and less on function-specific code offsets. A Jupyter Notebook was created that performs the actions listed in Table  II.
TABLE II: Preprocessing Steps
@@
Function names
that were not defined,
but were called
within the code
were given a
unique identifier
@_undefined_func_unique_name() extfuncvar_x ()
@@
labels, stack pointers,
and numeric metadata
were renamed
@@
!x
Digits were isolated
from the surrounding
characters and converted
into their word forms
123
one two three
@@
opcurl
dotmark
commark
  • a
    x is a number used to track the order in which it was defined.
@@
The first of these techniques, CBOW, generates a vector that uniquely represents a particular word using the words before and after. It is especially good at identifying similarities between words because similar words are often found within the same word patterns. For example, in the sentence “One of my favorite foods to eat is ice cream,” the word ice could easily be replaced with whipped or even sour. Thus, these words would have vectors that highlight that they are similar. However, the word road would have a vector that is quite different than ice, whipped, or sour because the words before and after the word are generally quite different from those for the other words. The CBOW technique predicts a missing word using the words surrounding the target word. The second technique used by word2vec, Skip-Gram, attempts to predict surrounding words using a target word, unlike CBOW, which predicts a target word based on the surrounding words.
The word2vec: CBOW and word2vec: Skip-Gram models were created using the gensim (4.3.0) Python library. Each preprocessed LLVM function was tokenized using the word tokenizer function of the nltk (3.8.1) library. The model was built using all the 48,157 samples from the embedding model dataset. There were 332 unique tokens used to train these models, but the vector size of the output embedding was set to one hundred. The window size was set to five, and the minimum count required for each token was set to one.
@@ II-E BERT and RoBERTa Model Creation
The code for the BERT model was released in 2019 [devlin2019bert] and stands for Bidirectional Encoder Representations from Transformers. BERT is a bidirectional encoder that learns token embeddings by masking out tokens within the input data and attempting to predict the missing token. Whereas some of the current models at the time were unidirectional, BERT proposed a bidirectional model that scored substantially higher on various NLP tasks than its predecessors. RoBERTa (Robustly Optimized BERT pre-training Approach) was also released in 2019 [liu2019roberta]. RoBERTa is similar to BERT, with some optimizations used to increase its performance.
In this research, the custom BERT model for learning semantics between LLVM instructions was built using the BertForMaskedLM class from Hugging Face’s transformers library (4.43.3), while the BertWordPieceTokenizer class from the tokenizers library (0.19.1) was used to tokenize LLVM functions. Similarly, the RoBERTa model was built using the RobertaForMaskedLM and ByteLevelBPETokenizerclass classes from these libraries. In both cases, 12 hidden layers with 100 nodes were used, each with an output layer of 100. The maximum vocabulary length was 2048, and batch sizes were limited to one due to GPU memory limitations. Training the BERT and RoBERTa models requires a training and validation dataset. Thus, the 48,157 samples from the embedding model dataset were split into two datasets; the training dataset comprised 90% of the overall dataset, and the validation dataset was 10% of the dataset.
For both the BERT and RoBERTa embedding models, the loss of training and validation steps was shown every 1,000 steps. The best loss in the validation set for the BERT and RoBERTa models was achieved after step 54,000 as shown in Table III. These were the embedding models used to create the embeddings provided to the LSTM models.
TABLE III: BERT/RoBERTa Five Best Loss Scores by Validation Step
BERT RoBERTa
Step Training Validation Training Validation
50,000 2.555300 2.496149 2.558400 2.494601
51,000 2.555500 2.497498 2.557400 2.488815
52,000 2.551600 2.493027 2.559900 2.485950
53,000 2.546700 2.495847 2.548000 2.487699
54,000 2.554400 2.492988 2.557500 2.483047
@@ II-F LSTM Neural Network Configuration
For the current research project, the LSTM neural networks comprised two hidden layers with 128 neurons each and 50 epochs to train each network. The Leaky Rectified Linear Unit (LeakyReLU) activation function was chosen because recurrent neural networks are prone to the problem of vanishing or exploding gradients [RN25, p. 335], and the LeakyReLU activation function helps avoid this problem. Each layer of the LSTM model was created using the TensorFlow and Keras libraries. A dropout rate of 20% was used between the LSTM layers, and a single dense layer of a single neuron was used in the output layer. The dataset was randomized and divided into two subsets. The first subset was used to train each neural network and consisted of 80% (i.e., 3,041 samples) of the overall dataset; the second subset was used for model validation and consisted of 20% (i.e., 761) of the total dataset. The same training and validation subsets were used for each neural network.
All neural networks were trained twice: once with the hidden layers of the embedding model frozen and once with them unfrozen. Each LSTM was created using several different hyperparameters to better understand the effects of optimizers and learning rates on model performance. The hyperparameters used for training each neural network are as follows:
  • Adam optimizer: learning rates of 0.01, 0.001, and 0.0001
  • SGD optimizer: a learning rate and momentum of 0.01
  • SGD: a learning rate of 0.0001 and momentums of 0.01, 0.001, and 0.0001.
Figure 2: Training Neural Networks to Detect Buffer Overflows
III Results 111 The training of several LSTM models with BERT embeddings, along with two models using RoBERTa embeddings, is nearing completion. A new preprint with final results will be submitted to arXiv in the coming weeks.
Accuracy, precision, recall, and F1 score were evaluated for each LSTM neural network created from the embeddings generated from each embedding model. Although each neural network was trained with and without freezing the embedding model’s hidden layers, only the results from the unfrozen models are included due to their superior performance. The metric tables in the following section do not include models with hyperparameters that did not achieve accuracy gains beyond the base accuracy of placing all samples in a single category.
III-A word2vec: Skip-Gram Metrics
@@
As shown in Table IV, the model with the highest performance using word2vec Skip-Gram embeddings achieved an accuracy of 92.0%, making it the model with the best performance in this study. This model was generated using the Adam optimizer with a learning rate of 0.001. The SGD optimizer results with a learning rate of 0.0001 and a momentum of 0.01, 0.001, and 0.0001 struggled to achieve accuracy gains beyond the base accuracy of placing all samples in a single category. However, these models consistently achieved lower loss values for each epoch, which indicates that additional epochs may eventually increase model performance.
TABLE IV: word2vec: Skip-Gram Metrics
@@
Hyper- parameters Epoch Loss Accuracy Precision Recall F1- Score
SGD -LR: 0.01 -Mom: 0.01 45 0.3663 80.9% 68.2% 90.0% 77.6%
Adam -LR: 0.01 35 0.4127 81.1% 69.1% 87.5% 77.2%
Adam -LR: 0.001 43 0.1848 92.0% 85.2% 94.6% 89.6%
Adam -LR: 0.0001 48 0.3282 84.0% 72.1% 91.8% 80.8%
III-B word2vec: CBOW Metrics
The metrics in Table  V were recorded when each model reached its highest accuracy. The highest accuracy, 87.5%, was obtained when the Adam optimizer was used with a learning rate of 0.001. LSTMs trained with the SGD optimizer and a low learning rate exhibited poor performance. However, similar to models trained with Skip-Gram embeddings, consistent improvements in loss scores across epochs suggest that additional training could have further enhanced performance.
TABLE V: word2vec: CBOW Metrics
@@
Hyper- parameters Epoch Loss Accuracy Precision Recall F1- Score
SGD -LR: 0.01 -Mom: 0.01 43 0.3417 83.0% 71.8% 88.5% 79.3%
SGD -LR: 0.0001 -Mom: 0.01 50 0.5920 69.8% 67.9% 33.3% 44.7%
SGD -LR: 0.0001 -Mom: 0.001 42 0.6246 69.0% 64.2% 34.8% 45.1%
SGD -LR: 0.0001 -Mom: 0.0001 37 0.6159 68.7% 63.4% 34.8% 44.9%
Adam -LR: 0.01 46 0.3531 83.4% 70.7% 93.5% 80.6%
Adam -LR: 0.001 34 0.2528 87.5% 78.0% 91.8% 84.3%
Adam -LR: 0.0001 45 0.2990 86.5% 77.0% 90.0% 83.0%
III-C RoBERTa Metrics